The line integral of the magnetic field on a closed path surrounding a wire has the value 10.0μT·m. What is the current in the wire?


1
Expert's answer
2021-03-08T08:43:49-0500

According to Ampere's law,

"\oint B.dl=\mu _0 I"

The current passing through the wire is

"I= \frac{\oint B.dl}{ \mu _0}
 = \frac{10.0 \times 10^{-6}T·m}{4 \pi \times 10^{-7} H/m}=7.96 A"